# HCT-QA: A Benchmark for Question Answering on Human-Centric Tables

### v2

@@ 4.1 Experimental setup
Large Models – 100B+ (Closed)
gpt-o1-mini 0.67 0.665 0.617 0.741 0.742 0.708 0.775 0.756 0.716 0.74 0.733 0.699 - - 0.691 0.685
gpt-4o 0.627 0.606 0.515 0.691 0.677 0.6 0.601 0.591 0.535 0.646 0.63 0.565 0.521 0.439 0.635 0.616
gpt-35 0.412 0.415 0.315 0.443 0.447 0.356 0.375 0.395 0.333 0.45 0.468 0.386 0.391 0.309 0.419 0.425
Average (Large) 0.57 0.562 0.482 0.625 0.622 0.555 0.584 0.581 0.528 0.612 0.61 0.55 0.304 0.249 0.582 0.575
Medium Models – 27B-70B (Open)
Qwen2.5-72B-Instruct 0.574 0.582 0.479 0.622 0.629 0.541 0.505 0.519 0.465 0.594 0.595 0.52 0.502 0.491 0.579 0.586
Llama-3.1-70B-Instruct 0.489 0.474 0.374 0.531 0.515 0.413 0.476 0.473 0.411 0.469 0.454 0.386 0.451 0.427 0.49 0.475
Llama-3.3-70B-Instruct 0.467 0.494 0.401 0.531 0.555 0.475 0.456 0.489 0.431 0.453 0.476 0.415 0.402 0.436 0.471 0.497
gemma-2-27b-it 0.473 0.466 0.356 0.536 0.525 0.415 0.396 0.406 0.351 0.479 0.464 0.379 0.378 0.322 0.477 0.469
gemma-3-27b-it 0.441 0.424 0.338 0.5 0.479 0.404 0.419 0.417 0.359 0.46 0.44 0.377 0.34 0.297 0.448 0.431
Mixtral-8x7B-Instruct-v0.1 0.279 0.278 0.197 0.307 0.307 0.229 0.248 0.274 0.227 0.306 0.323 0.259 0.17 0.159 0.284 0.286
Average (Medium) 0.454 0.453 0.358 0.505 0.502 0.413 0.417 0.43 0.374 0.46 0.459 0.389 0.374 0.355 0.458 0.457
Small Models – 3B-12B (Open)
Meta-Llama-3-8B-Instruct 0.309 0.285 0.181 0.324 0.293 0.201 0.237 0.245 0.191 0.269 0.255 0.184 0.235 0.181 0.302 0.28
Llama-3.1-8B-Instruct 0.223 0.215 0.138 0.235 0.214 0.127 0.204 0.204 0.165 0.248 0.237 0.177 0.153 0.132 0.227 0.218
Qwen1.5-7B-Chat 0.181 0.168 0.14 0.188 0.172 0.145 0.156 0.157 0.14 0.194 0.184 0.169 0.138 0.104 0.183 0.17
Qwen2-7B-Instruct 0.24 0.297 0.208 0.274 0.309 0.214 0.227 0.288 0.243 0.235 0.302 0.251 0.278 0.266 0.242 0.299
Qwen2.5-7B-Instruct 0.35 0.38 0.264 0.396 0.429 0.311 0.27 0.336 0.284 0.337 0.373 0.291 0.257 0.242 0.35 0.382
gemma-2-9b-it 0.377 0.353 0.243 0.409 0.368 0.254 0.287 0.299 0.253 0.347 0.345 0.265 0.349 0.275 0.373 0.351
gemma-3-12b-it 0.421 0.372 0.261 0.471 0.41 0.297 0.355 0.352 0.295 0.386 0.352 0.28 0.425 0.29 0.418 0.372
Mistral-7B-Instruct-v0.3 0.208 0.201 0.151 0.216 0.221 0.174 0.188 0.205 0.176 0.215 0.222 0.185 0.144 0.115 0.209 0.206
Mathstral-7B-v0.1 0.274 0.263 0.192 0.301 0.284 0.213 0.266 0.268 0.233 0.29 0.278 0.237 0.211 0.19 0.278 0.267
FanarC 0.27 0.243 0.178 0.238 0.21 0.151 0.271 0.26 0.235 0.273 0.258 0.215 0.179 0.145 0.267 0.243
aya-23-8B 0.113 0.098 0.069 0.111 0.095 0.063 0.11 0.101 0.077 0.134 0.135 0.111 0.068 0.058 0.115 0.103
Phi-3.5-mini-instruct 0.114 0.102 0.078 0.137 0.118 0.092 0.081 0.076 0.062 0.161 0.145 0.115 0.162 0.125 0.121 0.108
Average (Small) 0.257 0.248 0.175 0.275 0.26 0.187 0.221 0.233 0.196 0.257 0.257 0.207 0.195 0.157 0.257 0.25
VISION-TEXT VLMs/MLMs
Large Models – 100B+ (Closed)
gpt-4o-vision 0.486 0.615 0.513 0.583 0.704 0.613 0.31 0.396 0.328 0.571 0.674 0.597 - - 0.487 0.623
Small Models – 3B-12B (Open)
Pixtral-12B-2409 - 0.435 0.312 - 0.532 0.405 - 0.4 0.344 - 0.547 0.465 - - - 0.459
InternVL2-4B - 0.29 0.193 - 0.372 0.262 - 0.254 0.202 - 0.385 0.298 - - - 0.309
InternVL2_5-8B-MPO - 0.239 0.152 - 0.302 0.203 - 0.228 0.181 - 0.376 0.298 - - - 0.264
Phi-3-vision-128k-instruct - 0.408 0.3 - 0.445 0.338 - 0.273 0.23 - 0.442 0.357 - - - 0.411
Phi-3.5-vision-instruct - 0.244 0.159 - 0.365 0.262 - 0.178 0.14 - 0.375 0.29 - - - 0.272
Idefics3-8B-Llama3 - 0.382 0.269 - 0.458 0.337 - 0.351 0.284 - 0.438 0.354 - - - 0.396
Qwen2-VL-7B-Instruct - 0.06 0.045 - 0.066 0.057 - 0.071 0.052 - 0.045 0.038 - - - 0.059
llava-1.5-7b-hf - 0.094 0.068 - 0.109 0.095 - 0.085 0.062 - 0.067 0.056 - - - 0.091
llava-v1.6-mistral-7b-hf - 0.139 0.092 - 0.174 0.119 - 0.105 0.088 - 0.15 0.113 - - - 0.143
llava-v1.6-vicuna-7b-hf - 0.158 0.106 - 0.204 0.136 - 0.132 0.098 - 0.169 0.125 - - - 0.163
Molmo-7B-D-0924 - 0.376 0.276 - 0.472 0.368 - 0.282 0.225 - 0.465 0.383 - - - 0.394
paligemma2-10b-ft-docci-448 - 0.102 0.066 - 0.111 0.081 - 0.093 0.062 - 0.065 0.049 - - - 0.098
Average (Small) - 0.244 0.17 - 0.301 0.222 - 0.204 0.164 - 0.294 0.235 - - - 0.255
4.2 Analysis and findings
(RQ1) How do these models perform overall on HCT-QA?
@@ 4.2 Analysis and findings
As we can see in Table 1, the latest versions of large closed-weight models like gpt-o1-mini (text modality) perform reasonably well, achieving 69% Precision and 68% Recall. This shows that even for state-of-the-art models, there is ample room for improvement on this task. Furthermore, Qwen2.5-72B-Instruct, a considerably smaller model, achieves a Recall of 58%, which is only 3 points less than the Recall of gpt-4o. This showcases the potential for medium-sized models in this task. On average, large models achieve a Recall of 57% compared to 45% for medium-sized models, showing that the performance gap between models of medium and large sizes is still small. However, with an average Recall of 25%, small models lag behind their larger versions. Small vision models with an average Recall of 25.5% are on par with small text-only models. But Pixtral-12B VLM outperforms the best Qwen2.5-7B LLM of comparable size by 7 points in Recall, emphasizing the potential of vision models in capturing the structural complexity of HCTs that may be missed in a textual representation (check Appendix D.5 for details). Overall, there remains clear room for improvement for both LLMs and VLMs.
(RQ2) Are models improving over time?
Newer models are improving their reasoning capabilities and thus their performance on HCT-QA. The Qwen family of models shows continuous improvement as Qwen1.5-7B, Qwen2-7B, and Qwen2.5-7B respectively achieve recall of 17%, 29%, and 38%. The same trend can be seen with the gpt and gemma families of models (Table 1). It should be noted that the Llama-3 family is an exception to this trend; Llama-3-8B (28% Recall) performs 6 points better than its later iteration Llama-3.1-8B (21.8% Recall). On the VLM side, Phi models also display this trend of decreasing performance for newer iterations (Phi-3-vision has 41% Recall while Phi3.5-vision reaches only 27% Recall).
@@
Table 2: Model Precision & Recall across different HCTs input formats. Confidence intervals given in Appendix D.7.
Model Nonsensical-CSV Regular-CSV HTML
Precision Recall Precision Recall Precision Recall
gpt-4o 0.378 0.264 0.476 0.407 0.559 0.491
Llama-3.3-70B-Instruct 0.216 0.226 0.38 0.407 0.456 0.497
gemma-3-27b-it 0.211 0.17 0.344 0.305 0.457 0.429
gemma-3-12b-it 0.218 0.136 0.417 0.31 0.483 0.399
Qwen2.5-7B-Instruct 0.257 0.183 0.339 0.285 0.431 0.38
Llama-3.1-8B-Instruct 0.062 0.051 0.155 0.139 0.274 0.241
Meta-Llama-3-8B-Instruct 0.207 0.149 0.243 0.193 0.322 0.28
Mathstral-7B-v0.1 0.16 0.13 0.2 0.188 0.271 0.276
Average 0.214 0.164 0.319 0.279 0.366 0.339
Table 3: Fine-tuned model performances on real HCTs. Confidence intervals given in Appendix D.7
Model Fine-tuning Set Precision Recall CC Score
No FT 0.304 $|$ 0.262 0.279 $|$ 0.234 0.183 $|$ 0.171
Llama-3-8B $|$ FanarC-9B Real 0.600 $|$ 0.638 0.617 $|$ 0.646 0.525 $|$ 0.552
Real + Synth. 0.841 $|$ 0.864 0.849 $|$ 0.867 0.706 $|$ 0.734
(RQ3) Does the input format of the HCT matter in text-modality?
For the text models, we tested the effect of representing the table in HTML format instead of CSV. We used 1,000 synthetic HCTs (15,000 QA pairs) using 8 text-modality LLMs of different sizes. The results are shown in Table 2. We can see that the HTML format shows an improvement of 6% compared to canonical Regular-CSV, which is intuitive since HTML is a more expressive format that can better retain the complex structure of the HCTs.
@@
We also explore the importance of having semantically coherent and meaningful wording within the HCTs. We created Nonsensical-CSV versions of the Regular-CSV tables where we replace all the words like ”Meat”, ”2024”, or ”Gender” in the HCTs and their corresponding QA pairs, with nonsensical terms like ”AAjw”, ”CCab”, and ”DDnq”. Although the complexity of the questions and table structure remains unchanged, we see that the semantics drastically impact the performance, with a decrease of Recall by 12 points on Nonsensical-CSV compared to Regular-CSV. Even large models like gpt-4o cannot compensate for the lack of semantics as its Recall score drops by 14 points. This is in stark contrast to SQL-based query engines where semantics have no effect whatsoever. This emphasizes the importance of having clear and semantically coherent names and values in HCTs when using LLMs for QA tasks.
(RQ4) Does fine-tuning models for this task improve performance?
We fully fine-tuned the FanarC-9B and Meta-Llama-3-8B-Instruct models on two SFT sets: (i) Real HCTs which has 7,867 (80% of total) HCT-QA triplets on Real World HCTs only, and (ii) Real + Synthetic HCTs which has 7,867 triplets on Real World HCTs and 62,000 triplets on Synthetic HCTs (total 69,867). The test set contains 992 triplets on Real World HCTs only. We ensure no table in the test set is seen during training (Details in Appendix D.6). The results are shown in Table 3. Fine-tuning drastically increases the performance of both models. Llama-3-8B Recall increases by 34 points when fine-tuned on Real HCTs only, and by 57 points when fine-tuned on Real + Synthetic HCTs compared to its off-the-shelf (No FT) Recall. FanarC-9B Recall increases by 41 points and 63 points, respectively, for each fine-tuning set. These results showcase the potential of models to learn HCT-QA tasks given enough training data. Furthermore, it reflects the importance of our synthetic data generator since including them more than tripled all scores on Real World HCTs.
Table 4: Effect of table properties on Recall. Only properties with significant effect ($p<0.05$) are shown

### v3

@@ V-A Experimental setup
Large Models – 100B+ (Closed Weight)
ChatGPT 3.5 175B 0.413 0.315 0.445 0.356 0.385 0.333 0.459 0.386 0.345 0.279 0.409 0.334
ChatGPT 4o Unknown 0.663 0.571 0.720 0.639 0.650 0.583 0.719 0.654 0.563 0.365 0.663 0.562
ChatGPT 4.1 Unknown 0.621 0.545 0.640 0.576 0.650 0.600 0.667 0.605 0.543 0.415 0.624 0.548
Average - 0.566 0.477 0.602 0.524 0.562 0.505 0.615 0.548 0.484 0.353 0.566 0.481
Medium Models – 27B–72B (Open Weight)
Llama-3.1-70B-Instruct 70B 0.586 0.484 0.639 0.551 0.585 0.519 0.632 0.555 0.512 0.323 0.591 0.486
Llama-3.3-70B-Instruct 70B 0.573 0.497 0.636 0.571 0.548 0.530 0.645 0.595 0.499 0.361 0.580 0.511
Qwen2.5-72B-Instruct 72B 0.621 0.532 0.673 0.593 0.589 0.540 0.678 0.612 0.582 0.339 0.629 0.523
Qwen3-32B 32B 0.523 0.435 0.578 0.491 0.538 0.479 0.576 0.503 0.458 0.227 0.535 0.427
gemma-2-27b-it 27B 0.549 0.444 0.617 0.514 0.509 0.466 0.631 0.552 0.433 0.207 0.548 0.437
gemma-3-27b-it 27B 0.535 0.445 0.598 0.514 0.527 0.469 0.622 0.560 0.476 0.275 0.552 0.453
Average (Medium) - 0.565 0.473 0.624 0.539 0.550 0.500 0.631 0.563 0.493 0.289 0.572 0.473
Small Models – 4B–20B (Open Weight)
gemma-2-9b-it 9B 0.475 0.360 0.537 0.413 0.409 0.357 0.497 0.422 0.383 0.171 0.460 0.345
gemma-3-12b-it 12B 0.498 0.371 0.545 0.412 0.465 0.401 0.543 0.443 0.449 0.195 0.500 0.364
Qwen1.5-7B-Chat 7B 0.220 0.173 0.268 0.220 0.194 0.182 0.314 0.268 0.124 0.059 0.224 0.180
Qwen2-7B-Instruct 7B 0.333 0.273 0.356 0.305 0.283 0.289 0.375 0.341 0.295 0.107 0.328 0.263
Qwen2.5-7B-Instruct 7B 0.417 0.343 0.454 0.384 0.361 0.383 0.483 0.445 0.334 0.124 0.410 0.336
Qwen3-14B 14B 0.490 0.430 0.535 0.482 0.458 0.454 0.577 0.547 0.394 0.156 0.491 0.414
Mistral-7B-Instruct-v0.1 7B 0.071 0.062 0.077 0.065 0.032 0.027 0.143 0.129 0.062 0.040 0.077 0.065
Mistral-7B-Instruct-v0.2 7B 0.119 0.108 0.120 0.123 0.137 0.124 0.169 0.173 0.059 0.035 0.121 0.112
Mistral-7B-Instruct-v0.3 7B 0.295 0.219 0.321 0.255 0.253 0.222 0.342 0.298 0.151 0.068 0.273 0.213
Mathstral-7B-v0.1 7B 0.326 0.266 0.387 0.320 0.184 0.169 0.362 0.329 0.215 0.097 0.295 0.236
Meta-Llama-3-8B-Instruct 8B 0.344 0.241 0.374 0.268 0.334 0.302 0.376 0.301 0.262 0.100 0.338 0.242
Llama-3.1-8B-Instruct 8B 0.363 0.257 0.393 0.278 0.327 0.290 0.416 0.342 0.310 0.119 0.362 0.257
Phi-3-mini-4k-instruct 4B 0.233 0.167 0.265 0.185 0.223 0.198 0.339 0.267 0.211 0.107 0.254 0.185
Phi-3.5-vision-instruct 4B 0.274 0.193 0.235 0.170 0.190 0.143 0.245 0.189 0.108 0.059 0.211 0.151
Phi-4-mini-instruct 4B 0.225 0.156 0.131 0.086 0.044 0.037 0.154 0.116 0.062 0.021 0.123 0.083
Fanar-C-9B-Instruct 9B 0.431 0.322 0.463 0.351 0.473 0.416 0.522 0.451 0.437 0.217 0.465 0.351
Average (Small) - 0.320 0.246 0.341 0.270 0.273 0.249 0.366 0.316 0.241 0.105 0.308 0.237
VISION-TEXT VLMs
Large Models – 100B+ (Closed Weight)
ChatGPT 4o Vision Unknown 0.543 0.513 0.638 0.613 0.348 0.328 0.618 0.597 - - 0.537 0.513
Small Models – 3B–12B (Open Weight)
Pixtral-12B-2409 12B 0.378 0.289 0.471 0.372 0.300 0.273 0.457 0.411 0.364 0.161 0.394 0.301
InternVL2-4B 4B 0.236 0.181 0.281 0.199 0.148 0.162 0.305 0.297 0.265 0.102 0.247 0.188
InternVL2_5-8B-MPO 8B 0.193 0.129 0.270 0.196 0.177 0.170 0.349 0.288 0.208 0.076 0.240 0.172
Phi-3-vision-128k-instruct 4B 0.214 0.187 0.280 0.252 0.210 0.199 0.339 0.317 0.241 0.135 0.257 0.218
Phi-3.5-vision-instruct 4B 0.208 0.138 0.314 0.227 0.148 0.120 0.332 0.269 0.265 0.133 0.253 0.177
llava-1.5-7b-hf 7B 0.067 0.061 0.096 0.092 0.067 0.057 0.042 0.039 0.030 0.006 0.061 0.051
llava-v1.6-vicuna-7b-hf 7B 0.123 0.091 0.157 0.118 0.095 0.077 0.133 0.107 0.115 0.032 0.125 0.085
Molmo-7B-D-0924 7B 0.130 0.067 0.164 0.100 0.092 0.075 0.147 0.103 0.156 0.027 0.138 0.074
Average (Small) - 0.194 0.143 0.254 0.194 0.155 0.142 0.263 0.229 0.205 0.084 0.214 0.158
Note: ChatGPT-4o Vision was not run on synthetic HCT images due to API constraints
V-B Analysis and findings
@@
(RQ1) How do language models perform overall on HCT-QA? As shown in Table IV, the relatively recent versions of large closed-weight LLMs like ChatGPT 4o and ChatGPT 4.1 on the text modality perform reasonably well, achieving F1 scores of 66% and 62% respectively. This shows that even for state-of-the-art LLMs, there is ample room for improvement on this task. Despite ChatGPT 4o being the best performing model overall, we see that its predecessor ChatGPT 3.5 scores an average F1 of 40% which is less than all of the medium sized open-weight LLMs. This highlights the progress more recent LLMs have made in their reasoning abilities despite their much smaller size. Furthermore, Qwen2.5-72B-Instruct achieves an average F1 of 62.9%, which is only 3.4 percentage points less than ChatGPT 4o’s score, despite its size of 72B parameters being considerably smaller than the OpenAI model. Similarly, Qwen2.5-72B-Instruct’s CC Score of 52% is only 4 percentage points less that ChatGPT 4o CC Score of 56%. On average, large LLMs achieve an F1 of 56% compared to 57% for medium-sized LLMs. Even, when excluding the lagging ChatGPT 3.5, large LLMs have an average F1 score of 64% which is only 7 percentage points higher than the medium-sized LLMs’ average. This showcases the potential for medium-sized LLMs on this task.
With an average F1 of 30%, small LLMs lag behind their larger competitors. Even among small LLMs, gemma-3-12b-it’s average F1 score of 50% outperforms ChatGPT 3.5 by nearly 10 percentage points. However, we see a considerable drop in gemma-3-12b-it’s CC Score; only 36.4% which is far below the average for medium and large models. This indicates that it does get a lot of the questions partially correct but struggles to provide the fully correct answer. Moreover, small VLMs with an average F1 of 21.4% are far worse off than small LLMs. But, among the VLMs, Pixtral-12B outperforms many LLMs such as Llama-3.1-8B-Instruct, Mistral-7B-Instruct-v0.3, and Qwen2-7B-Instruct. This indicates that VLMs do have the potential to compete with similarly sized LLMs on HCT-QA, possibly by utilizing the visual features of HCTs that may be lost in a textual representation. Overall, there remains clear room for improvement for both LLMs and VLMs. Furthermore, the F1 scores of models on synthetic HCTs is, on average, slightly less than on real-world HCTs. This shows that our synthetic generator produces HCTs and QA pairs that are challenging for all models. However, there is a considerable drop in the CC Scores on synthetic HCTs. This could be due to many of the synthetic QA pairs requiring answers with many more values than the QA pairs on real-world HCTs–as the number of values in the answer increases, the models’ ability to generate a completely correct answer significantly degrades. For example, the Qwen2.5-72B-Instruct model under the synthetic HCT tables has the highest F1 score while its CC score is relatively low.
In Figure 8 we also compare how models perform on QA pairs created using our three methods. Synthetic questions appear to be the most challenging which is expected as the majority require Row Filtering, Expression on Columns, and Aggregation together. On average, models of all sizes perform best on the manually-generated QA pairs. This indicates that our model-generated and synthetic questions provide a slightly harder challenge for the evaluated models.
(RQ2) Are models improving over time? There appears to be no consistent trend with regards to newer generations of models outperforming their predecessors. The Qwen family of small LLMs shows continuous improvement as Qwen1.5-7B, Qwen2-7B, Qwen2.5-7B, and Qwen3-14B all outperform their predecessor with average F1 scores of 22.4%, 32.8%, 41%, and 49.1% respectively. This shows that the Qwen family is improving the performance of their small LLMs by an average of 9% per generation. The same can be seen with their CC Scores as Qwen3-14B is 23 percentage points better than Qwen1.5-7B-Chat. This shows that the models are getting much better at giving the complete correct answer. We see similar trends with the Mistral 7B LLMs which also improve by nearly 9% per generation and Gemma LLMs that improve by 4%. Conversely, the Phi family of small LLMs shows consistent deterioration in each successive generation with Phi-3-mini-4k-instruct achieving an average F1 score of 25.4% which is nearly double that of Phi-4-mini-instruct, despite the latter being released 8 months later. The ChatGPT family shows a mixed trend with a significant improvement of 26 percentage points between ChatGPT 3.5 and ChatGPT 4.0, but then ChatGPT 4.1 under-performs its predecessor by 4 percentage points. There are also cases like gemma-2-27b-it and gemma-3-27b-it which shows the newer generation improving by a mere 0.4 percentage points in terms of average F1. We also see a similar trend with small VLMs as Phi-3-vision-128k-instruct outperforms its successor, Phi-3.5-vision-instruct, by 0.4 percentage points. Thus, we cannot say that newer versions of models outperform their older versions across the board.
Figure 8: Model performance by size across different QA creation types
@@
TABLE V: Model F1 scores across different HCTs input formats
c Model CSV Markdown HTML Text Obfuscated HTML ChatGPT 4.1 0.484 0.529 0.543 0.499 ChatGPT 4o 0.541 0.567 0.563 0.531 Qwen2.5-72B-Instruct 0.422 0.453 0.582 0.508 Llama-3.3-70B-Instruct 0.431 0.443 0.499 0.398 Qwen2.5-7B-Instruct 0.301 0.311 0.334 0.358 Llama-3.1-8B-Instruct 0.191 0.262 0.311 0.277 Average 0.395 0.426 0.472 0.42
TABLE VI: Finetuned model performances across different test sets.
Model Name Finetuning Set Test Sets (F1 Score)
Real HCTs Synthetic HCTs HiTab [13]
Llama-3.1-8B-Instruct None 0.351 0.333 0.538
Llama-3.1-8B-Instruct Real HCTs 0.593 0.534 0.606
Llama-3.1-8B-Instruct Synth. HCTs 0.471 0.879 0.561
Llama-3.1-8B-Instruct Real & Synth. HCTs 0.595 0.886 0.616
(RQ3) Does the input format of the HCT matter in text-modality? For the text modality LLMs, we tested the effect of representing the HCT in CSV and Markdown formats instead of HTML. The results are shown in Table V. The HTML format outperforms CSV and Markdown formats by an average F1 score difference of 7.7 and 4.6 percentage points, respectively. This result is intuitive as HTML is a more expressive format that can better retain the complex structure of the HCTs. Some models are more sensitive than others to the table format. Qwen2.5-72B-Instruct performs 13 percentage points better with the HTML format than with Markdown and shows a similar difference between the HTML and CSV formats as well. Whereas, ChatGPT 4o shows a relative indifference to the table format as there is only a 0.2 percentage point difference in F1 scores between the Markdown and HTML formats, with the former performing minutely better. Our results show that while large LLMs like ChatGPT can perform well with both Markdown and HTML, medium and small sized LLMs perform much better with an HTML representation of the HCT.
@@
We also explore the effect of semantically coherent and meaningful text within the HCTs. We created Text Obfuscated-HTML versions of the HCT HTMLs where we (i) replace all column names with ambiguous terms (ii) replace all text cell values in the HCTs with similar ambiguous terms, and (iii) replace the corresponding QA pairs with the respective ambiguous terms from the obfuscated table. Although the complexity of the questions and table structure remains unchanged, we see that the text obfuscation negatively impacts model performance. LLMs on average, experience a decrease in F1 score of 5 percentage points on Text Obfuscated-HTML format when compared to the original HTMLs. Even a large model like ChatGPT 4.1 cannot compensate for the lack of semantics as its F1 score drops by 4.4 percentage points.
(RQ4) Does finetuning models improve performance? We finetuned the Llama-3.1-8B-Instruct model on three FT sets: (i) Real HCTs which has 7,867 (80% of total) HCT-QA triplets on real world HCTs only, (ii) Synth. HCTs which has 54,186 (80% of total) HCT-QA triplets on synthetic HCTs only, (iii) Real + Synth. HCTs which combines both the  Real HCTs and  Synth. HCTs sets (62,053 in total). We then evaluate the models on the respective Real HCTs and Synth. HCTs test sets. We ensure that no HCT in the test set is seen during finetuning. Furthermore, to test the generalization of finetuning in HCT-QA we test on the HiTab dataset [13]. The results are shown in Table VI. Finetuning on the Real + Synth. HCTs set drastically increases the performance of the model across the board. The F1 score improvement, compared to the non-finetuned model is 24.1 percentage points on real HCTs from HCT-QA, 55 percentage points on synthetic HCTs from HCT-QA, and nearly 8 percentage points on the out of domain HiTab dataset. Furthermore, finetuning on the Synth. HCTs set yields an F1 score improvement of 12 percentage points on the real HCTs test set. This reinforces the usefulness of our synthetic data generator since training only on synthetic HCTs generalizes well to real-world HCTs.
TABLE VII: Effect of table properties on F1. Only properties with significant effect ($p<0.05$) are shown
@@
(a) Refer to caption
(b) Refer to caption
Figure 9: F1 scores for models across a subset of question properties (a) and aggregation types (b). Models are ordered by decreasing size from top to bottom, and by decreasing average scores from left to right.